A firm with a 10 percent cost of capital is considering a project for this year’s capital budget. The project’s expected after-tax cash flows are as follows:

Year:

0

1

2

3

4

Cash flow:

-$10,000

$3,200

$3,400

$4,700

$4,700


Calculate the project’s net present value (NPV).

a.

$6,000.00

b.

$2,460.35

c.

$2,236.68

d.

$4,545.45

e.

$928.22
Answer #1

Time Cash Flows Present Value I Present Value B = (1/1.10)^n C = A*B WNO -10,000.00 3,200.00 3,400.00 4,700.00 4,700.00 0.909

So, Net Present value = $2,460.35

Option 'B' is correct.